# Probing cluster magnetism with embedded and background radio sources in Planck clusters

### v1

@@ 3.2 Comparison to models
In the simple scenario of random magnetic field orientations in cells of size $\Lambda_{c}$ kpc, which have some magnetic field strength and electron density, the observed RM is the result of a random walk process. Because of the central limit theorem, the distribution of RMs is then expected to be a Gaussian distribution with zero mean, and variance given by (e.g. Murgia et al., 2004)
@@
In reality, the magnetic field structure will more closely resemble a random field with fluctuations on many spatial scales, and both the magnetic field strength and electron density will scale with radius. Thus comparing observations to more realistic scenarios requires simulated magnetic fields. We followed the approach explained in Section 4 of Osinga et al. (2022) to generate mock rotation measure and depolarisation images for all clusters in our sample which have X-ray observations available (99/124). An example of a mock RM and depolarisation image is shown in Figure 1 for a random cluster in our sample. In these models, the magnetic field is assumed to be a three-dimensional Gaussian random field with a single-power law spectrum characterised by the following parameters: $B_{0}$, $\eta$, $n$, $\Lambda_{\mathrm{min}}$ and $\Lambda_{\mathrm{max}}$. The first two denote the variables that parameterise the magnetic field, assumed to follow (e.g. Bonafede et al., 2010)
@@
between minimum and maximum fluctuations scales that are denoted by $\Lambda_{\mathrm{min}}$ and $\Lambda_{\mathrm{max}}$ in image space, respectively. In the picture of Kolmogorov turbulence, $n=11/3$, but this was found observationally to take values between $n=1$ and $n=4$ (Murgia et al., 2004; Govoni et al., 2006; Guidetti et al., 2008; Bonafede et al., 2010; Vacca et al., 2010, 2012; Govoni et al., 2017; Stuardi et al., 2021; Osinga et al., 2022). We have computed all models on $1024^{3}$ pixel grids, to simulate all clusters homogenously. Here, one pixel represents 3 kpc, and clusters are thus simulated out to about $1.5R_{\mathrm{500}}$, with a minimum fluctuation scale of $\Lambda_{\mathrm{min}}=6$ kpc. These models will be compared to observations in various ways, as detailed in the next section.
@@ 4.1 Average magnetic field strength
In Figure 2, we plot the RRM as a function of projected distance to the nearest cluster centre (denoted as $a$ throughout this work). A clear trend is visible, with the scatter in RRM decreasing with a larger distance to the cluster centre. We determined the corrected RRM scatter for all sources within $1R_{\mathrm{500}}$, using Eq. 6, and find that the RRM scatter of cluster members and background sources are similar [^3] All uncertainties on $\sigma_{\mathrm{RRM}}$ are calculated by 1 000 bootstraps., being $200\pm 33$ and $219\pm 66$ rad m-2, respectively, as shown in Table 2. In reality, for a given magnetic field strength, the RRM scatter of background sources should be $\sqrt{2}$ times that of cluster members, as cluster members are on average located at the mid-plane of the cluster. However, cluster members are found preferentially at smaller radii, where the scatter in RM is larger due to generally larger magnetic field strength and electron densities.
Table 2: Corrected standard deviation of RM in rad m-2, as defined in Eq. 6 for different subsets and projected radii. We refer to Fig. 6 for the corrected IQR as a function of radius.
@@
444
The mean value of the RMs agrees with zero as a function of radius, as shown in Figure 3, consistent with random magnetic field orientations along the line of sight. Assuming the simple random walk scenario denoted by Equation 7, we find that the most rudimentary estimate of the line-of-sight magnetic field strength is given by (e.g. Böhringer et al., 2016)
$\left(\frac{B_{\mathrm{||}}}{\mu\,G}\right)=2.46\left(\frac{\sigma_{\mathrm{RM}}/200}{\,\mathrm{rad\,m}^{-2}}\right)\left(\frac{n_{e}/10^{-3}}{\,\mathrm{cm}^{-3}}\right)\\ \left(\frac{\Lambda_{c}/10}{\,\mathrm{kpc}}\right)^{-1/2}\left(\frac{L/1000}{\,\mathrm{kpc}}\right)^{-1/2}$ (10)
@@
To sample a more physical property, we plot in Figure 4 the observed rotation measures as a function of ICM electron column density. This plot is less populated, as we now only show sources that are detected within a projected radius $a<2R_{\mathrm{500}}$, where the Chandra-derived column density values are reliable. The scatter in rotation measure significantly increases with increasing column density, and the preferential sampling of cluster members at high column densities (i.e. low radii) is clearly pronounced. Following Böhringer et al. (2016), we calculated the scatter in rotation measure in bins of column density, from which the average magnetic field strength along the line of sight can be calculated as
@@ 4.4 Constraining the magnetic field power spectrum
The median observed RM-DP ratio across all sources was found to be $113\pm 12$ rad m-2, with the uncertainty calculated by 1 00010001\,0001 000 bootstraps, with background sources and cluster members having similar values of $106\pm 17$ rad m-2 and $119\pm 14$ rad m-2, respectively. Figure 10 shows the median observed RM ratio compared with simulated values for sources sampled at similar positions in the simulated RM and depolarisation images. It is clear from the figure that models with most of the magnetic field energy on small scales (i.e. $n<3$) cannot reproduce the observed RM ratio, mainly because those models result in too low values of $|\mathrm{RM}|$ due to the rapidly fluctuating magnetic field along the line of sight. Instead, models with $n\geq 3$ provide a good fit for various values of the maximum correlation scale $\Lambda_{\mathrm{max}}$. Lowering $\Lambda_{\mathrm{max}}$ has an analogous effect to lowering $n$, namely decreasing the coherence length of the magnetic field along the line of sight. This thus results in a smaller average $|\mathrm{RM}|$ while the effect on depolarisation is less significant, as this is measured on scales below the observing beam (less than typically 15 kpc, although dependent on cluster redshift). Thus, the observed data best matches $n\geq 3$.
@@ 5 Discussion
Other statistical studies probing many unresolved sightlines such as Böhringer et al. (2016); Stasyszyn & de los Rios (2019) have also not observed a strong increase of $\sigma_{\mathrm{RRM}}$ at low radii, implying the same observational bias. However, at higher frequencies, RM values as high as 10,000 rad m-2 have been observed in the centre of some clusters (e.g. Taylor & Perley, 1993). Higher frequency data that suffers less from depolarisation is thus needed to determine the value of $\eta$ accurately. This will be provided by, for example, the VLA Sky Survey (VLASS; Lacy et al., 2020) Depolarisation does not suffer as strongly from this observational bias at low radii, as upper limits on depolarisation fractions can still be set on sources that are significantly depolarised, given that they are detected at high signal-to-noise in total intensity. However, the depolarisation curves that we observed also do not agree with any of the Gaussian random field models explored in this paper, being generally flatter than models, as also seen in (Osinga et al., 2022).
Thus, we finally fit both the $\sigma_{\mathrm{RRM}}$ and depolarisation radial profile jointly by means of forward modelling the NCC clusters, since the CC clusters showed anomalous behaviour as explained below. None of the models could explain the RM and depolarisation data consistently, with observations finding a flatter profile than expected from the models. However, our models assume clusters are spherical over-densities without any surrounding large-scale structure, which is unrealistic since clusters will on average lie between one or two filaments (e.g. Pimbblet et al., 2004). The densities and possibly magnetic field strengths in the regions connecting the clusters to the filaments will be higher than our spherical $\beta$-like model, plausibly flattening the observed RM and depolarisation profiles. Comparisons with cosmological magneto-hydrodynamical (MHD) simulations are an obvious next step, where clusters are embedded in the large-scale structure around them (e.g. Marinacci et al., 2018).
@@
Finally, we checked for qualitative differences between merging and relaxed clusters. During cluster mergers, turbulence is injected on large scales, which can amplify magnetic fields, and drive large fluctuation scales (e.g. Vacca et al., 2018). Relaxed clusters, on the other hand, are expected to have smaller fluctuation scales, as the energy injected by a previous merger has dissipated through cascades to smaller and smaller scales. We found a significant difference between the merging and relaxed cluster samples. Merging clusters show higher values of $\sigma_{\mathrm{RRM}}$ in the region $0.5<r/R_{\mathrm{500}}$ than relaxed clusters, and sources inside relaxed clusters probe a flat $\sigma_{\mathrm{RRM}}$ profile as a function of radius. Higher RM scatter in merging clusters was also observed by Stasyszyn & de los Rios (2019), although with larger radial bin sizes. If a scaling between $B$ and $n_{e}$ is assumed, then CC clusters should show a steeper radial profile near the centre, as CC clusters often have higher central electron densities, as also shown in Fig 21. It is unclear what is causing the relative flatness of the $\sigma_{\mathrm{RRM}}$ profile in the sample of cluster members inside CC clusters, so CC clusters were not used for the full RM and depolarisation forward model.
@@ 6 Conclusion
    Jointly modelling both the depolarisation and rotation measure of sources in a forward modelling approach, we find that the observations cannot be explained by the simple model of a Gaussian random magnetic field that follows $B\propto n_{e}^{\eta}$. This implies that more advanced modelling of cluster magnetic fields is required.
In this work, we implicitly assumed that all clusters have the same magnetic field parameters, while in reality, this might be a function of dynamical state, mass, or redshift. The universality of cluster magnetic fields has not been thoroughly tested (e.g. Govoni et al., 2017) and might significantly affect our results. Future observations with more sensitive telescopes such as MeerKAT and the Square Kilometre Array could test this assumption by detecting enough polarised sightlines through single clusters such that stacking is not required.
@@ Acknowledgements.
EO and RJvW acknowledge support from the VIDI research programme with project number 639.042.729, which is financed by the Netherlands Organisation for Scientific Research (NWO). EO thanks both Wout Goesaert and Joppe Swart for the careful looks at his code and Bryan Gaensler for providing comments on the manuscript. The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. Basic research in radio astronomy at the U.S. Naval Research Laboratory is supported by 6.1 Base Funding. This paper has made use of observational material taken with an NRAO instrument. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID 2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID 2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID 2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. The Legacy Surveys project is honoured to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University This research has made use of the NASA/IPAC extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. This research has made use of NASA’s Astrophysics Data System (ADS). This research made use of Astropy, a community-developed core Python package for Astronomy. This research has made use of ChatGPT to restructure text and code.

### v2

@@ 3.2 Comparison to models
In the simple scenario of random magnetic field directions in cells of size $\Lambda_{c}$ kpc, which have some uniform magnetic field strength and uniform electron density, the observed RM is the result of a random walk process. Because of the central limit theorem, the distribution of RMs is then expected to be a Gaussian distribution with zero mean, and variance given by (e.g. Murgia et al., 2004)
@@
In reality, the magnetic field structure will more closely resemble a random field with fluctuations on many spatial scales, and both the magnetic field strength and electron density will scale with radius. Thus comparing observations to more realistic scenarios requires simulated magnetic fields. We followed the approach explained in Section 4 of Osinga et al. (2022) to generate mock rotation measure and depolarisation images for all clusters in our sample which have X-ray observations available (99/124). We assumed that the clusters have spherical and smooth electron density profiles, with best-fits to the de-projected electron density as a function of radius taken from Andrade-Santos et al. (2017), which are shown in Fig 21. In reality, there are likely density fluctuations in the ICM. However, these are difficult to measure and on the order of only a few per cent (e.g. Schuecker et al., 2004; Churazov et al., 2012; Sanders & Fabian, 2012). Magnetic field fluctuations are expected to be stronger. Therefore, we have modelled the magnetic field as a Gaussian random field.
An example of a mock RM and depolarisation image is shown in Figure 1 for a random cluster in our sample. To compare the mock images to observations, we sample a mock source for every observed source at the same projected radius. If the source was classified as a background source, the integration is done over the full grid, while for a cluster member, the integration is done assuming the source lies at the midplane of the cluster.
In these models, the magnetic field is assumed to be a three-dimensional Gaussian random field with a single-power law spectrum characterised by the following parameters: $B_{0}$, $\eta$, $n$, $\Lambda_{\mathrm{min}}$ and $\Lambda_{\mathrm{max}}$. The first two denote the variables that parameterise the magnetic field strength, assumed to follow
@@
The last three parameters encode the power spectrum of the magnetic field:
@@
[…]
@@ 4.1 Average magnetic field strength
In Figure 2, we plot the RRM as a function of projected distance to the nearest cluster centre (denoted as $a$ throughout this work). A clear trend is visible, with the scatter in RRM decreasing with a larger distance to the cluster centre. We determined the corrected RRM scatter for all sources within $1R_{\mathrm{500}}$, using Eq. 6, and find that the RRM scatter of cluster members and background sources are similar [^5] All uncertainties on $\sigma_{\mathrm{RRM}}$ are calculated by 1 000 bootstraps., being $200\pm 33$ and $219\pm 66$ rad m-2, respectively, as shown in Table 2. Assuming that clusters are spherically symmetrical on average, or at least have no bias towards our direction, and that cluster members are on average located at the mid-plane of the cluster, we expect the RRM scatter of background sources to be at least $\sqrt{2}$ times that of cluster members at the same projected radius. However, cluster members are found preferentially at smaller radii (e.g. Osinga et al., 2022, Fig. B1), where the scatter in RM is larger due to generally larger magnetic field strength and electron densities (e.g. Bonafede et al., 2010).
Table 2: Corrected standard deviation of RM in rad m-2, as defined in Eq. 6 for different subsets and projected radii.
@@
 [^6] We refer to Fig. 6 for the corrected IQR as a function of radius.
The mean value of the RMs agrees with zero as a function of radius, as shown in Figure 3, consistent with random magnetic field orientations along the line of sight. Assuming the simple random walk scenario denoted by Equation 7, we find that the most rudimentary estimate of the line-of-sight magnetic field strength is given by
$\left(\frac{B_{\mathrm{||}}}{\mu\,G}\right)=2.46\left(\frac{\sigma_{\mathrm{RM}}/200}{\,\mathrm{rad\,m}^{-2}}\right)\left(\frac{n_{e}/10^{-3}}{\,\mathrm{cm}^{-3}}\right)^{-1}\\ \left(\frac{\Lambda_{c}/10}{\,\mathrm{kpc}}\right)^{-1/2}\left(\frac{L/1000}{\,\mathrm{kpc}}\right)^{-1/2}$ (10)
@@
As we have information on the radial shape of the electron densities, we plot in Figure 4 the observed rotation measures as a function of ICM electron column density. This plot is less populated, as we now only show sources that are detected within a projected radius $a<2R_{\mathrm{500}}$, where the Chandra-derived column density values are reliable. The scatter in rotation measure significantly increases with increasing column density, and the preferential sampling of cluster members at high column densities (i.e. low radii) is clearly pronounced. Following Böhringer et al. (2016), we calculated the scatter in rotation measure in bins of column density, from which the average magnetic field strength along the line of sight can be calculated as
@@ 4.4 Constraining the magnetic field power spectrum
The median observed ratio of ${|\mathrm{RM}|}/(1-\mathrm{DP})$ across all sources was found to be $113\pm 12$ rad m-2, with the uncertainty calculated by 1 00010001\,0001 000 bootstraps, with background sources and cluster members having similar values of $106\pm 17$ rad m-2 and $119\pm 14$ rad m-2, respectively. Figure 10 shows the median ratio compared with simulated values for sources sampled at similar positions in the simulated RM and depolarisation images. It is clear from the figure that models with most of the magnetic field energy on small scales (i.e. $n<3$) cannot reproduce the observed RM ratio, mainly because those models result in too low values of $|\mathrm{RM}|$ due to the rapidly fluctuating magnetic field along the line of sight. Instead, models with $n\geq 3$ provide a good fit for various values of the maximum correlation scale $\Lambda_{\mathrm{max}}$. Lowering $\Lambda_{\mathrm{max}}$ has an analogous effect to lowering $n$, namely decreasing the coherence length of the magnetic field along the line of sight. This thus results in a smaller average $|\mathrm{RM}|$ while the effect on depolarisation is less significant, as this is measured on scales below the observing beam (less than typically 15 kpc, although dependent on cluster redshift). Thus, the observed data best matches $n\geq 3$.
@@ 5 Discussion
Other statistical studies probing many unresolved sightlines such as Böhringer et al. (2016); Stasyszyn & de los Rios (2019) have also not observed a strong increase of $\sigma_{\mathrm{RRM}}$ at low radii, implying the same observational bias. However, at higher frequencies, RM values as high as 10,000 rad m-2 have been observed in the centre of some clusters [^8] Depending on the electron density and magnetic field structure in the cluster, Taylor & Perley (1993) determined that a magnetic field strength of 30 $\mu$G was required in case of the Hydra A cluster. (e.g. Taylor & Perley, 1993; Baidoo et al., 2023). Higher frequency data that suffers less from depolarisation is thus needed to determine the value of $\eta$ accurately. This will be provided by, for example, the VLA Sky Survey (VLASS; Lacy et al., 2020) Depolarisation does not suffer as strongly from this observational bias at low radii, as upper limits on depolarisation fractions can still be set on sources that are significantly depolarised, given that they are detected at high signal-to-noise in total intensity. However, the depolarisation curves that we observed also do not agree with any of the Gaussian random field models explored in this paper, being generally flatter than models, as also seen in (Osinga et al., 2022).
Thus, we finally fit both the $\sigma_{\mathrm{RRM}}$ and depolarisation radial profile jointly by means of forward modelling the NCC clusters, since the CC clusters showed anomalous behaviour as explained below. None of the models could explain the RM and depolarisation data consistently, with observations finding a flatter profile than expected from the models. However, our models assume clusters are spherical over-densities without any surrounding large-scale structure, which is unrealistic since clusters will on average lie between one or two filaments (e.g. Pimbblet et al., 2004). The densities and possibly magnetic field strengths in the regions connecting the clusters to the filaments will be higher than our spherical $\beta$-like model, plausibly flattening the observed RM and depolarisation profiles. Additionally, our models do not account for any fluctuations in the electron density with a similar or different power spectrum than that of the magnetic field. Although there are indications from simulations that the magnetic field and thermal density power spectra could be similar at small scales (e.g. Domínguez-Fernández et al., 2019), this is difficult to constrain observationally and expected to be on the order of a few per cent (e.g. Churazov et al., 2012). Additional density fluctuations would generally increase the modelled RM scatter, and thus require slightly lower magnetic field strengths to fit observations. Finally, the assumption that the magnetic fields are Gaussian is in contrast with expectations if the fields are generated by the fluctuation dynamo mechanism, which predicts spatially intermittent fields with non-Gaussian statistics (e.g. Schekochihin et al., 2004; Vazza et al., 2018; Seta et al., 2020; Kopyev et al., 2022) Comparisons with cosmological magneto-hydrodynamical (MHD) simulations are an obvious next step, with realistic turbulent gas profiles, potential non-Gaussian magnetic fields and clusters that are embedded in the large-scale structure around them (e.g. Marinacci et al., 2018).
@@
Finally, we checked for qualitative differences between merging and relaxed clusters. During cluster mergers, turbulence is injected on large scales, which can amplify magnetic fields, and drive large fluctuation scales (e.g. Vacca et al., 2018). Relaxed clusters, on the other hand, are expected to have smaller fluctuation scales, as the energy injected by a previous merger has dissipated through cascades to smaller and smaller scales. We found a significant difference between the merging and relaxed cluster samples. Merging clusters show higher values of $\sigma_{\mathrm{RRM}}$ in the region $0.5<r/R_{\mathrm{500}}$ than relaxed clusters, and sources inside relaxed clusters probe a flat $\sigma_{\mathrm{RRM}}$ profile as a function of radius. Higher RM scatter in merging clusters was also observed by Stasyszyn & de los Rios (2019), although with larger radial bin sizes. If a scaling between $B$ and $n_{e}$ is assumed, then CC clusters should show a steeper radial profile near the centre, as CC clusters often have higher central electron densities, as also shown in Fig 21. It is unclear what is causing the relative flatness of the $\sigma_{\mathrm{RRM}}$ profile in the sample of cluster members inside CC clusters, so CC clusters were not used for the full RM and depolarisation forward model. In the future, exploring a two-component magnetic field model for CC clusters might alleviate the inconsistency, analogous to two-component models that are used to fit CC cluster thermal electron density profiles (e.g. Vikhlinin et al., 2006).
@@ 6 Conclusion
    Jointly modelling both the depolarisation and rotation measure of sources in a forward modelling approach, we find that the observations cannot be explained by the simple model of a Gaussian random magnetic field that follows $B\propto n_{e}^{\eta}$. Although $\eta=0$ is slightly preferred over $\eta=0.5$, more advanced modelling of cluster magnetic fields is required.
In this work, we implicitly assumed that all clusters have the same magnetic field parameters, while in reality, this might be a function of dynamical state, mass, or redshift. The universality of cluster magnetic fields has not been thoroughly tested (e.g. Govoni et al., 2017) and might significantly affect our results. Future observations with more sensitive telescopes such as MeerKAT and the Square Kilometre Array could test this assumption by detecting enough polarised sightlines through single clusters such that stacking is not required. Such observations should take into account that Equation 2 might not hold for all sources, as diffuse radio emission from clusters (e.g. van Weeren et al., 2019) might induce complexity into the Faraday spectra.
Data availability
Table 3 is available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
We thank the referee, Amit Seta, for a thorough report that has improved the quality of this manuscript. EO and RJvW acknowledge support from the VIDI research programme with project number 639.042.729, which is financed by the Netherlands Organisation for Scientific Research (NWO). RJvW acknowledges support from the ERC Starting Grant ClusterWeb 804208. EO thanks both Wout Goesaert and Joppe Swart for the careful looks at his code and Bryan Gaensler for providing comments on the manuscript. The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. Basic research in radio astronomy at the U.S. Naval Research Laboratory is supported by 6.1 Base Funding. This paper has made use of observational material taken with an NRAO instrument. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID 2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID 2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID 2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. The Legacy Surveys project is honoured to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University This research has made use of the NASA/IPAC extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. This research has made use of NASA’s Astrophysics Data System (ADS). This research made use of Astropy, a community-developed core Python package for Astronomy. This research has made use of ChatGPT to restructure text and code.